Calculate the density of a 7.0% solution of sodium chloride in water if it has a volume of 40.0 mL and a mass of 42.08 grams
1
Expert's answer
2016-05-25T08:43:02-0400
rho = m/V = 42.08 / 40 = 1.052 (g/ml)